OpenStudy (anonymous):

b^2-4b=0 solve medals will be given

OpenStudy (anonymous):

x(1)=0 and x(2)=4

OpenStudy (id21):

\[b^2-4b=0 \] \[b(b-4)=0\] b=0 or b-4=0; b=4

OpenStudy (anonymous):

thanks guys!